OpenStudy (anonymous):

Help!what is the least common denominator of 4/x^2y+5/y^2?

OpenStudy (anonymous):

i can give you a problem that is the same but not the same numbers @Brookebrookiee

OpenStudy (anonymous):

Okay, thank you, that works.

OpenStudy (anonymous):

is x2 - x - 2 & x2 - 3x + 2 ok

OpenStudy (anonymous):

Yes

OpenStudy (jinx.exe):

6x+5y/x2+4xy+4y2/x+2y/5 Final result : 20x3y + 30x3 + 2x2y + 20xy2 + 25y ————————————————————————————————— 5x2 Step by step solution : Step 1 : Raise y to the 2nd power Exponentiaion : Equation at the end of step 1 : y y2 2y (((6x+(5•————))+4xy)+(4•——))+—— (x2) x 5 Step 2 : Raise x to the 2nd power Exponentiaion : Equation at the end of step 2 : y 4y2 2y (((6x+(5•——))+4xy)+———)+—— x2 x 5 Step 3 : 5y Simplify 6x + —— x2 Rewriting the whole as an Equivalent Fraction : 3.1 Adding a fraction to a whole Rewrite the whole as a fraction using x2 as the denominator : 6x 6x • x2 6x = —— = ——————— 1 x2 Equivalent fraction : The fraction thus generated looks different but has the same value as the whole Common denominator : The equivalent fraction and the other fraction involved in the calculation share the same denominator Adding fractions that have a common denominator : 3.2 Adding up the two equivalent fractions Add the two equivalent fractions which now have a common denominator Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible: 6x • x2 + 5y 6x3 + 5y ———————————— = ———————— x2 x2 Equation at the end of step 3 : (6x3 + 5y) 4y2 2y ((—————————— + 4xy) + ———) + —— x2 x 5 Step 4 : 6x3+5y Simplify —————— + 4xy x2 Rewriting the whole as an Equivalent Fraction : 4.1 Adding a whole to a fraction Rewrite the whole as a fraction using x2 as the denominator : 4xy 4xy • x2 4xy = ——— = ———————— 1 x2 Trying to factor as a Sum of Cubes : 4.2 Factoring: 6x3 + 5y Theory : A sum of two perfect cubes, a3 + b3 can be factored into : (a+b) • (a2-ab+b2) Proof : (a+b) • (a2-ab+b2) = a3-a2b+ab2+ba2-b2a+b3 = a3+(a2b-ba2)+(ab2-b2a)+b3= a3+0+0+b3= a3+b3 Check : 6 is not a cube !! Ruling : Binomial can not be factored as the difference of two perfect cubes Adding fractions that have a common denominator : 4.3 Adding up the two equivalent fractions (6x3+5y) + 4xy • x2 4x3y + 6x3 + 5y ——————————————————— = ——————————————— x2 x2 Equation at the end of step 4 : (4x3y + 6x3 + 5y) 4y2 2y (————————————————— + ———) + —— x2 x 5 Step 5 : 4x3y+6x3+5y 4y2 Simplify ——————————— + ——— x2 x Trying to factor a multi variable polynomial : 5.1 Factoring 4x3y + 6x3 + 5y Try to factor this multi-variable trinomial using trial and error Factorization fails Calculating the Least Common Multiple : 5.2 Find the Least Common Multiple The left denominator is : x2 The right denominator is : x Number of times each Algebraic Factor appears in the factorization of: Algebraic Factor Left Denominator Right Denominator L.C.M = Max {Left,Right} x 2 1 2 Least Common Multiple: x2 Calculating Multipliers : 5.3 Calculate multipliers for the two fractions Denote the Least Common Multiple by L.C.M Denote the Left Multiplier by Left_M Denote the Right Multiplier by Right_M Denote the Left Deniminator by L_Deno Denote the Right Multiplier by R_Deno Left_M = L.C.M / L_Deno = 1 Right_M = L.C.M / R_Deno = x Making Equivalent Fractions : 5.4 Rewrite the two fractions into equivalent fractions Two fractions are called equivalent if they have the same numeric value. For example : 1/2 and 2/4 are equivalent, y/(y+1)2 and (y2+y)/(y+1)3 are equivalent as well. To calculate equivalent fraction , multiply the Numerator of each fraction, by its respective Multiplier. L. Mult. • L. Num. (4x3y+6x3+5y) —————————————————— = ————————————— L.C.M x2 R. Mult. • R. Num. 4y2 • x —————————————————— = ——————— L.C.M x2 Adding fractions that have a common denominator : 5.5 Adding up the two equivalent fractions (4x3y+6x3+5y) + 4y2 • x 4x3y + 6x3 + 4xy2 + 5y ——————————————————————— = —————————————————————— x2 x2 Equation at the end of step 5 : (4x3y + 6x3 + 4xy2 + 5y) 2y ———————————————————————— + —— x2 5 Step 6 : 4x3y+6x3+4xy2+5y 2y Simplify ———————————————— + —— x2 5 Checking for a perfect cube : 6.1 4x3y+6x3+4xy2+5y is not a perfect cube Trying to factor a multi variable polynomial : No factorization found Calculating the Least Common Multiple : 6.2 Find the Least Common Multiple The left denominator is : x2 The right denominator is : 5 Number of times each prime factor appears in the factorization of: Prime Factor Left Denominator Right Denominator L.C.M = Max {Left,Right} 5 0 1 1 Product of all Prime Factors 1 5 5 Number of times each Algebraic Factor appears in the factorization of: Algebraic Factor Left Denominator Right Denominator L.C.M = Max {Left,Right} x 2 0 2 Least Common Multiple: 5x2 Calculating Multipliers : 6.3 Calculate multipliers for the two fractions Denote the Least Common Multiple by L.C.M Denote the Left Multiplier by Left_M Denote the Right Multiplier by Right_M Denote the Left Deniminator by L_Deno Denote the Right Multiplier by R_Deno Left_M = L.C.M / L_Deno = 5 Right_M = L.C.M / R_Deno = x2 Making Equivalent Fractions : 6.4 Rewrite the two fractions into equivalent fractions L. Mult. • L. Num. (4x3y+6x3+4xy2+5y) • 5 —————————————————— = —————————————————————— L.C.M 5x2 R. Mult. • R. Num. 2y • x2 —————————————————— = ——————— L.C.M 5x2 Adding fractions that have a common denominator : 6.5 Adding up the two equivalent fractions (4x3y+6x3+4xy2+5y) • 5 + 2y • x2 20x3y + 30x3 + 2x2y + 20xy2 + 25y ———————————————————————————————— = ————————————————————————————————— 5x2 5x2 Final result : 20x3y + 30x3 + 2x2y + 20xy2 + 25y ————————————————————————————————— 5x2

OpenStudy (anonymous):

well nevermind haha @Brookebrookiee

OpenStudy (anonymous):

I still don't understand how I solve it?

OpenStudy (anonymous):

@Jinx.exe will my answer be x+5?

OpenStudy (jinx.exe):

look at the final result

OpenStudy (anonymous):

It says 5x2?

OpenStudy (jinx.exe):

it does